# ToMAP: Training Opponent-Aware LLM Persuaders with Theory of Mind

### v1

@@ Appendix E Details about Persuasion Strategies
In Section 5.3, we analyze the persuasion strategies used in ToMAP. Table 8 provides An Example of each strategy, and we also provide the prompt used to annotate strategies with GPT-4o. In addition, we show a typical case of ToMAP’s arguments, where the model effectively utilizes multiple persuasion strategies to convince its opponent.
@@
An Example of ToMAP’s Argument CLAIM: Productivity is overvalued and prioritizing it above all else can be detrimental to well-being and creativity. ARGUMENT: While Bob’s arguments are valid, it is crucial to consider the context in which productivity is pursued. The current economy and workplace environments often prioritize productivity to make as much profit as possible [Preemptive Rebuttal: consider why Bob emphasizes productivity and refute accordingly.]. However, the balance between productivity and other crucial aspects of life is critical. By highlighting the ways in which excessive productivity can lead to stress, burnout, and a lack of creativity [Evidential Appeal: list bad impacts of over-emphasizing productivity to support Alice’s claim.], we argue that the current paradigm of prioritizing productivity above all else is indeed detrimental. This shows that the pursuit of productivity must be balanced with other important factors [Concession: first aim for a balance instead of directly refuting Bob’s view.] to ensure that individuals can achieve true personal achievement and contribute effectively to the economy without sacrificing their well-being and creative capabilities.
@@
Appendix F The initial attitude of persuaders
The prompt design in Appendix B makes sure the persuadee disagrees with the target claims initially, leaving ample room for persuasion. Specifically, the persuadees’ average initial agreement scores on the three benchmarks are shown in Table 9, which confirms our prompt effectively set the original attitudes.
Table 9: Initial scores for each persuadee across datasets.
Persuadee CMV Anthropic args.me
QWen-7B 25.29 33.00 31.66
LLaMa-8B 20.45 18.00 23.66
Phi-4 19.92 24.33 20.53

### v2

@@ Appendix F Details about Persuasion Strategies
In Section˜5.4, we analyze the persuasion strategies used in ToMAP. Table˜10 provides An Example of each strategy, and we also provide the prompt used to annotate strategies with GPT-4o. In addition, we show a typical case of ToMAP’s arguments, where the model effectively utilizes multiple persuasion strategies to convince its opponent.
An Example of ToMAP’s Argument CLAIM: Prioritizing productivity above all else can be detrimental to well-being and creativity. ARGUMENT: While Bob’s arguments are valid, it is crucial to consider the context in which productivity is pursued. The current economy and workplace environments often prioritize productivity to make as much profit as possible [Preemptive Rebuttal: consider why Bob emphasizes productivity and refute accordingly.]. However, the balance between productivity and other crucial aspects of life is critical. By highlighting the ways in which excessive productivity can lead to stress, burnout, and a lack of creativity [Evidential Appeal: list bad impacts of over-emphasizing productivity to support Alice’s claim.], we argue that the current paradigm of prioritizing productivity above all else is indeed detrimental. This shows that the pursuit of productivity must be balanced with other important factors [Concession: first aim for a balance instead of directly refuting Bob’s view.] to ensure that individuals can achieve true personal achievement and contribute effectively to the economy without sacrificing their well-being and creative capabilities.